1. You are thinking of buying an apartment for $20,000. The apartment will be worth 545,000 in 1 years. Compare with the choice that you put the money in the bank (1) Should you buy the apartment if r- 0.05? (2) Should you buy it if r-0.10
Answer #1

In bank the deposit is always calculated on quarterly basis

n = 10 years

Formula to calculate Amount that is deposited in bank that is compounded quarterly

A = amount = P ( 1+r/4)(4n)

P = 20000

I) r = 0.05

A = 20000 (1+(0.05)/4)(4*10)

A = 20000*1.6436

A = 32872 $

Yes it is better to buy the apartment as at r =0.05 we will get an amount of $32872 which is less than $45000

2) r = 0.1

A = 20000 (1+(0.1/4))(4*10)

A = 20000* 2.685

A= $53700

It is better to keep that amount in bank at r=0.1 as we get an amount of $53700 which is greater than $45000